Suppose you need to have $52,557.00 in an account 19.00 years from today and that the account pays 8.00%. How much do you have to deposit into the account 6.00 years from today?

Answer Format: Currency: Round to: 2 decimal places.
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

52557=P*(1.08)^13

P=52557/1.08^13

=52557*0.367697924

=$19325.10(Approx).